# An Adaptive Gradient Regularization Method

### v2

@@ 3.3 AGR Formulation
$\alpha_{i,j}=\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}$ (2)
@@ Appendix A Appendix
$\begin{split}\left|\left|\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})\right|\right|^{2}_{2}=&\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}\leq\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}\\ \end{split}$ (5)
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})=&\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})\\ \end{split}$ (6)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=-\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ \end{split}$ (7)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}-2\nabla_{\textbf{w}_{i,j}\mathcal{L}}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (8)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ \end{split}$ (9)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1+\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}+2\nabla_{\textbf{w}_{i,j}}\mathcal{L}\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{\left|\sum\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (10)

### v3

@@ 3.3 AGR Formulation
$\alpha_{i,j}=\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}$ (2)
@@ Appendix A Appendix
$\begin{split}\left|\left|\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})\right|\right|^{2}_{2}=&\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}\leq\left|\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|\right|^{2}_{2}\\ \end{split}$ (5)
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})=&\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})\\ \end{split}$ (6)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=-\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ \end{split}$ (7)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+-\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}-2\nabla_{\textbf{w}_{i,j}\mathcal{L}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (8)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}(1-\frac{\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})&=\nabla_{\textbf{w}_{i,j}}\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|-\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ \end{split}$ (9)
@@
$\begin{split}\nabla_{\textbf{w}_{i,j}}\Psi(\nabla_{\textbf{w}_{i,j}}\mathcal{L})&=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}(1+\frac{\nabla_{\textbf{w}_{i,j}}\mathcal{L}}{\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|})^{2}+\frac{\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{w}_{i,j}}\mathcal{L}\nabla_{\textbf{w}_{i,j}}^{2}\mathcal{L}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}+2\nabla_{\textbf{w}_{i,j}}\mathcal{L}\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+(\nabla{\textbf{w}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &=\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\frac{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|+\nabla_{\textbf{W}_{i,j}}\mathcal{L})^{2}}{(\sum\left|\nabla_{\textbf{w}_{i,j}}\mathcal{L}\right|)^{2}}\\ &\leq\nabla^{2}_{\textbf{w}_{i,j}}\mathcal{L}\end{split}$ (10)